# Object-centric proto-symbolic behavioural reasoning from pixels

### v1

@@ 1 Introduction
A one-year old infant, before language is expressed [1], learns the efferent-afferent patterns of sensory stimulation and motor commands to a neural representation of the environment. From the sensorium, to abstract thought, and back to the minutiae of the sensorimotor domain, this interaction defines the feats of higher-order cognition and are the backdrop of the different levels of abstraction at which intelligent agents must operate, when they interact with the environment, which is not fully observable [2]. How sub-symbolic computations transform into higher-level cognitive representations of structures is far from being understood [3]. However, there is common accepted idea that suggest that there are intermediate representations that bridge cognitive reasoning and behaviour [4]. Hence, in the design of artificial agents, a key challenge is first to find a representational space that provides an effective interface between these disparate demands and second a mechanism that makes proper use of this representation to interact with the environment and produce behaviour. This is, we define reasoning from a behavioral perspective where the output of a reasoning process is always an action that physically interacts with the environment.
@@ 2 Related work, challenges and contribution
Structured representation learning appeared as a powerful way to introduce inductive biases and scale artificial intelligence to high-level cognition [8] and exploit symmetries and invariances that can be leveraged by decomposing scenes into their natural constituents [9]. Particularly, object-centric representations provide a natural interface between bottom-up (e.g., emergentist, connectionist) and top-down (e.g., graph based, probabilistic) approaches [7]. We summarize the closed-related works that influenced the proposed architecture classified in four topics [^1] Some of the works may have overlapping features in other topics.: scene understanding (e.g., visual segmentation), physics-informed and interaction (e.g., simulation), behaviour (e.g., robotics) and reasoning (e.g., object relations). We analyse these works from the prism of behavioural reasoning.
@@
Visual segmentation methods have reached maturity with high accuracy on segmentation in both static images [10] and videos [11]. From IODINE [10] that used IAI, recent methods evolved to fully amortized slot-attention architectures [12] and diffusion approaches [13]. These methods, which can track 2D and 3D objects in cluttered scenes [14], capture object collisions and discover new objects [15], are designed for perception and not for reasoning or control the segmented objects.
Physics-informed and interaction
Physics methods have also shown a great potential in handling complex non-linear interactions. These are strongly influenced by the simulation approach to cognition [4]. For instance, interaction [16] and propagation [17] networks are able to understand complex scene dynamics with multiple objects and even control them to reach a desired known state. Unfortunately, these methods require full observation of the objects states. Approaches that can work directly on pixels make use of ground-truth segmentation masks [18] or Visual interaction networks [19]. While these approaches have some kind of physical “reasoning” a full-fledged architecture that performs 1st order logic reasoning from pixels and transforms to behaviour (control) is still not fully accomplished.
@@
These works focus on moving the objects meaningfully and usually have full access to the objects states (e.g. [20]). OP3 [21] is an outstanding exception. However, it has two key limitations. First, it is restricted to a discrete action space of picking up and placing objects, with no continuous control. Second, it has no ability to learn tasks—it can only plan actions towards objectives specified ad hoc by means of a goal image that shows exactly what the scene in question ought to look like. Novel approaches use ReFs to describe the 3D representation of the object and control is solved by RRTs in the latent prediction dynamics and Model Predictive Control (MPC) to execute the actions [22]. Still, ground truth objects mask are used for training and inference. Some methods have overcome this constraint through entity-based segmentation—as a simplification of object-centric representation that uses the center of the object as the location of the entity—and RL to overcome with the generation of meaningful behaviours [23]. Unfortunately, this restricts the shape of objects to “point mass” like. Besides, there are object-centric approaches that instead of focusing on the interactive behaviour between the agent and the objects, they focus on view point matching [24] and agent navigation [25].
@@
Leaving out the literature on visual scene understanding with language, reasoning with object-centric representations as neurosymbols is, while the most promising, the least investigated area [26]. There are just a few object-centric studies on visual reasoning in static images [27]. Furthermore, there are relevant works from Large Language Models research (e.g., [28]), but they do not adhere to the grounding paradigm proposed, where reasoning should appear as an emergent property of learning a world physical model [29]. For instance in [30] the reasoning capabilities are much more expressive than our proposed approach but they use pre-trained skill, which are conditioned on the language generated by the LLM. Conversely, our architecture learns the world dynamics and interaction through unsupervised learning, harnessing the construction of the proto-symbols while interacting. This does not prevent the possibility of connecting proposed architecture to a LLMs similarly to [31], but through the preference network, which is already grounded.
@@ 2.1 Current challenges and architecture decisions
There are two major challenges at the unsupervised learning with object-centric representations: working with complex naturalistic scene images [13] and the generation of meaningful physical actions (i.e., interacting through continuous control) derived from the reasoning process when the input is high-dimensional (i.e., image) and the environment is partially-observable. This work focus on the latter. We show that simple conditional reasoning problems, such as, “if there is a heart move squares to the left, otherwise move squares to the right” ($(A\to B)\land(\neg A\to C)$) is already unsolvable for non-object-centric representations SOTA algorithms—See Sec. 4. An underlying problem for both challenges is to achieve high performance in world state dynamic estimation (world model learning), but allowing adaptation and generalization. We designed the object-centric backbone architecture to incorporate adaptation, following the free energy principle theory of information brain processing [32, 33], where the brain estimates the world by continuously approximating its internal model, learn by experience, to the real world through approximate Bayesian computation. To implement it, we opted for an iterative variational inference approach inspired by the IODINE [10] architecture. Whilst iterative methods have shown less performance than amortized approaches (e.g., slot-attention architecture [12]), at least at perceptual tasks, it provides Bayesian filtering and smoothing, thus allowing better counterfactual reasoning. Finally, another relevant challenge, which is not addressed in this work, is to deal with complex non-linearities of physical contact, such as collisions. Physics-based simulation based approaches [18] and Deep Reinforcement Learning (RL) [23] is showing the best results but they still provide the ground truth object masks or entail oversimplified segmentation in the visual input respectively.
@@ 2.2 Contribution
This work provides a neural network architecture (Fig. 1) that can learn 1st order conditional behavioural reasoning where proto-symbols [^2] We use the term proto-symbols as in visual segmentation we use for proto-objects [34], thus describing candidates for symbols and objects. take the form of object-centic representations and perception, control and preferences (desired internal state of the agent) are dynamical process computed through approximate Bayesian inference, in an unsupervised learning scheme from pixels. Our approach draws inspiration on natural intelligence, where the agent function is to generate behaviours through reasoning conditioned on perceptual cues by optimizing a single quantity: the variational free energy (or evidence lower bound) [32]. To this end, we combine bottom-up and top-down object-centric approaches [35, 36] following the emergentist approach (with unsupervised learning from pixels) but allowing probabilistic inference on the structured representation. This permits the agent to learn: $i)$ what is an object (scene understanding), $ii)$ how to mentally manipulate it through object-centric rules (reasoning) and $iii)$ generate physical behavior (i.e., continuous control). The agent can only perceive through visual input (2D image projection) of a synthetic world composed of 2D or 3D objects (see Fig. 2a for schematic). It can interact with the environment by applying forces to locations in the image, which corresponds to forces in the 2D/3D world. This architecture allows the agent to learn complex conditional rules, such as $(A\to B\land C)\land(\neg A\to D\land E)$, perform logical composition $(A\to B)\land(A\to C)\vdash A\to(B\land C)$ by learning two different rules separately but then they emerge combined during execution, and XOR operations, such as $(A\lor B)\to C)\land(A\land B\to D)$. Once the agent learns the rules, reasoning transforms into a behavioural response that changes the environment towards its internal preference. Furthermore, the agent, using the proposed architecture, shows online adaptation to unexpected changes in its environment, is robust to mild violations of its world model, and generalize to the number of objects.
@@ 4.1 Environment
We developed, the Active dSprites, which is an “activated” version of the various multi-dSprites datasets that have been used in previous work on object-based visual inference (e.g. [10, 12]). Not only does it include (continuous) dynamics, but these dynamics can can be acted on by an agent (through continuous control actions that accelerate the objects). Thus, active dSprites is an interactive environment, rather than a dataset. We implemented two different scenarios the active dSprites environment. The first, objects in can be 2.5-D shapes (e.g., squares, ellipses and hearts)—they have no depth dimension of their own, but can occlude each other within the depth dimension of the image. The second objects are 3D shapes (e.g., boxes, half-torus, cones) that can move in a $10\times 10\times 10$ m cubic space. These objects are much more complex to learn than the 2D version, as they have lighting and shadows, producing color gradients that need to be encoded. To randomize the environment, when an active dSprites instance is intialized, object shapes, positions, sizes and colors are all sampled uniformly at random. Initial velocities are drawn from a Normal distribution. Shape colors are sampled at discrete intervals spanning the full range of RGB-colors, while a background color is drawn from a set of evenly spaced grayscale values between black and white. Shapes are presented in random depth order. In the 3D environment the lighting sources are fixed. Code for the active dSprites environment can be found at github.com/neuro-ai-robotics/OBR.
@@ 4.2 Problem complexity analysis through baselines comparison
Figure 3: Complexity analysis through baseline comparison. Oracle algorithms have access to the objects location and dynamics, and define the upper-bound performance. RL algorithms (PPO and SAC) are tested with full observability and partial observability (RGB image pixels) (a) Algorithms performance at each frame of execution for a learnt single conditional rule $(A\to B)\land(\neg A\to C)$. Colored lines are the Mean Squared Error (MSE) between the objects location and the idealized goal location for 100 tested instances after training. Error bars are the standard deviation. (b) Average performance of the algorithms at the final frame. The lower the better.
Results are shown in Fig. 3. The left panel shows the task accuracy (Minimum Squarred Error, MSE) of the current object locations and the desired state in projected image) [^7] Note that this error is over the ground truth goal as we can use the environment information to obtain the true image projection of the goal.. All methods with full observability can solve the task, having all of them similar performance. When restricting the input to RGB images (pixels) standard SAC or PPO cannot learn the conditional rule. We can see that the MSE increases along the time. OBR, while it is not as accurate as the oracle it can learn the rule. The right panel shows the final performance of the compared algorithms. Oracle algorithms shows the upperbound performance as it is constrained by the dynamics of the environment.
These results first show, in line with previous works [21], that manipulating objects in the environment cannot be solved through non object-centric approaches. And highlights an important weakness of current pixel-based RL baselines. Second, it shows that a basic reasoning rule that implies interaction with the environment is already enough complex. Thus, learning multiple rules and composition in this active dSprite synthetic environment is relevant.
@@ 4.2.1 Abstract behavioural reasoning with conditional rules
We evaluated the capacity of the architecture to learn proto-symbolic behavioural rules and its generalization to the number of objects. We evaluated both 2D and 3D scenarios. Figure 4 shows an OBR agent execution for an“IfHeart” conditional $(A\to B)\land(\neg A\to C)$ and the XOR rules, and two randomized scenarios for each rule. In Fig. 4a, the first scenario, there is no heart, and boxes should be moved to the upper-left and ellipses to the middle-left. The second one, there is a heart present, and boxes, ellipses and hearts should be moved to the left top, middle and down respectively. The agent is passive until the 5th frame. The environment row shows the visual input and the red arrows describe the forces applied. The imagined goal row, describes the output from the preference network at each frame. Note that the preference is in encoded only in the latent representation and here, for visualization purposes, we show its reconstruction using the decoder network. We let the agent to perform three frames of iterative inference until the proposed goal is shown. The true goal is the idealized solution of this “IfHeart” task. Analogously, in Fig. 4b, the OBR agent, which has learnt the Heart XOR Square task, solves two 2d active dSprites scenarios.
@@ 4.2.2 Adaptation to changes in the environment
We evaluated the capacity of the OBR agent to recover from unexpected changes in the environment. For that purpose, we generate instances of the active dSprites that perform an object substitution in the middle of the execution. For instance, a Heart is changed by a Square during a conditional rule execution. Figure 6a describes two instances of the adaptation environment and Fig. 6b the statistical performance over 100 randomized instances. The object substitution experiment test the perception module, which needs to adapt fast to estimate the new situation, the preference network, which has to adapt the imagined goal depending on the objects that appear in the scene, and the control module that should provide the correct continuous actions to correct for the previously generated object movements. The true goal describes the idealized desired objects location after object substitution. The task error shows a decrease of performance until the system is able to recover and generate the proper behaviour to fulfill the conditional rule.
Figure 6: Adaptation. (a) Two instances of the active dSprites solved by the OBR agent which has learnt the “IfHeart” conditional rule: ((((Heart $\to$ Squares to top-right $\land$ Ellipses to middle-right)(¬$\land$($\neg$) ∧ ( ¬ Heart $\to$ Squares to top-left $\land$ Ellipses to middle-left $\land$ Hearts to down-left)))). At 4th frame one of the objects (in this case the heart is substituted by a square, forcing not only to change the object-centric perception but the conditional behaviour, as the resulting reasoning depends on the appearance of the Heart. (b) Statistical performance evaluation of 100 instances of the object substitution experiment. Blue line is the MSE and the error bars are the standard deviation.

### v2

@@ 1 Introduction
A one-year old infant, before language is expressed [1], learns the efferent-afferent patterns of sensory stimulation and motor commands to a neural representation of the environment. The pathway from the sensorium, to abstract thought, and back to the minutiae of the sensorimotor domain defines the feats of higher-order cognition and structures the different levels of abstraction at which intelligent agents must operate, when they interact with the environment [2]. How sub-symbolic computations transform into higher-level cognitive representations of structures is far from being understood [3]. However, there is common accepted idea that suggest that there are intermediate representations that bridge cognitive reasoning and behaviour [4]. Hence, in the design of artificial agents, a key challenge is first to find a representational space that provides an effective interface between these disparate demands, as well as a mechanism that makes proper use of this representation to interact with the environment and produce behaviour. In this sense, we approach reasoning from a behavioral perspective where the output of a reasoning process is always an action that physically interacts with the environment.
@@ 2 Related work, challenges and contribution
Structured representation learning appeared as a powerful way to introduce inductive biases and scale artificial intelligence to high-level cognition [8] and exploit symmetries and invariances that can be leveraged by decomposing scenes into their natural constituents [9]. Particularly, object-centric representations provide a natural interface between bottom-up (e.g., emergentist, connectionist) and top-down (e.g., graph based, probabilistic) approaches [7]. We summarize the closely related works that influenced the proposed architecture classified in four topics [^1] Some of the works may have overlapping features in other topics.: scene understanding (e.g., visual segmentation), physics-informed and simulation-based methods, robotics and reasoning (e.g., object relations). We analyse these works from the prism of behavioural reasoning.
@@
Visual segmentation methods have reached maturity with high accuracy on segmentation in both static images [10] and videos [11]. From IODINE [10] that used iterative amortized inference, recent methods evolved to slot-attention architectures [12] and diffusion approaches [13]. These methods, which can track 2D and 3D objects in cluttered scenes [14], capture object collisions and discover new objects [15], are designed for perception and not for reasoning about, or controlling the segmented objects.
Physics-informed and simulation-based methods
These approaches have also shown a great potential in handling complex non-linear interactions. These are strongly influenced by the simulation approach to cognition [4]. For instance, interaction [16] and propagation [17] networks are able to understand complex scene dynamics with multiple objects and even control them to reach a desired known state. Unfortunately, these methods require full observation of the objects states. Approaches that can work directly on pixels make use of ground-truth segmentation masks [18] or Visual interaction networks [19]. While these approaches have some kind of physical “reasoning” a full-fledged architecture that performs 1st order logic reasoning from pixels and transforms to behaviour (control) is still not fully accomplished.
@@
These works focus on moving the objects meaningfully and usually have full access to the objects states (e.g. [20]). OP3 [21] is an outstanding exception. However, it has two key limitations. First, it is restricted to a discrete action space of picking up and placing objects, with no continuous control. Second, it has no ability to learn tasks—it can only plan actions towards objectives specified ad hoc by means of a goal image that shows exactly what the scene in question ought to look like. Novel approaches use ReFs to describe the 3D representation of the object and control is solved by RRTs in the latent prediction dynamics and Model Predictive Control (MPC) to execute the actions [22]. Still, ground truth objects mask are used for training and inference. Some methods have overcome this constraint through entity-based segmentation—as a simplification of object-centric representation that uses the center of the object as the location of the entity—and RL to overcome with the generation of meaningful behaviours [23]. Unfortunately, this restricts the shape of objects to “point mass” like. Besides, there are object-centric approaches that, instead of focusing on the interactive behaviour between the agent and the objects, focus on view-point matching [24] and agent navigation [25].
@@
Leaving out the literature on visual scene understanding with language, reasoning with object-centric representations as neurosymbols is, while the most promising, the least investigated area [26]. There are just a few object-centric studies on visual reasoning in static images [27]. Furthermore, there are relevant works from Large Language Models research (e.g., [28]), but they do not adhere to the grounding paradigm proposed, where reasoning should appear as an emergent property of learning a world physical model [29]. For instance in [30] the reasoning capabilities are much more expressive than our proposed approach but they use pre-trained behaviours, which are conditioned on the language generated by the LLM. Conversely, our architecture learns the world dynamics and interaction through unsupervised learning, harnessing the construction of the proto-symbols while interacting. This does not prevent the possibility of connecting proposed architecture to a LLMs similarly to [31], but through the preference network, which is already grounded.
@@ 2.1 Current challenges and architecture decisions
There are two major challenges at the unsupervised learning with object-centric representations: working with complex naturalistic scene images [13] and the generation of meaningful physical actions (i.e., interacting through continuous control) derived from the reasoning process when the input is high-dimensional (i.e., image) and the environment is partially-observable. This work focus on the latter. We show that simple conditional reasoning problems, such as, “if there is a heart move squares to the left, otherwise move squares to the right” ($(A\to B)\land(\neg A\to C)$) is already unsolvable for non-object-centric representations SOTA algorithms—see Sec. 4. An underlying problem for both challenges is to achieve high performance in world state dynamic estimation (world model learning), but allowing adaptation and generalization. We designed the object-centric backbone architecture to incorporate adaptation, following the free energy principle theory of neural information processing [32, 33], where the brain estimates the world by continuously approximating its internal model, learned by experience, to the real world through approximate Bayesian computation. To implement it, we opted for an iterative variational inference approach inspired by the IODINE [10] architecture. Whilst iterative methods have shown less performance than amortized approaches (e.g., slot-attention architecture [12]), at least at perceptual tasks, it provides Bayesian filtering and smoothing, thus allowing better counterfactual reasoning. Finally, another relevant challenge, which is not addressed in this work, is to deal with complex non-linearities of physical contact, such as collisions. Physics-based simulation based approaches [18] and Deep Reinforcement Learning (RL) [23] are showing the best results but they still provide the ground truth object masks or entail oversimplified segmentation in the visual input respectively.
@@ 2.2 Contribution
This work provides a neural network architecture (Fig. 1) that can learn 1st order conditional behavioural reasoning where proto-symbols [^2] We use the term proto-symbols as in visual segmentation we use for proto-objects [34], thus describing candidates for symbols and objects. take the form of object-centic representations and perception, control and preferences (desired internal state of the agent) are dynamical processes computed through approximate Bayesian inference, in an unsupervised learning scheme from pixels. Our approach draws inspiration from natural intelligence, where the agent’s function is to generate behaviours through reasoning conditioned on perceptual cues by optimizing a single quantity: the variational free energy (or evidence lower bound) [32]. To this end, we combine bottom-up and top-down object-centric approaches [35, 36] following the emergentist approach (with unsupervised learning from pixels) but allowing probabilistic inference on the structured representation. This permits the agent to learn: $i)$ what is an object (scene understanding), $ii)$ how to mentally manipulate it through object-centric rules (reasoning) and $iii)$ generate physical behavior (i.e., continuous control). The agent can only perceive through visual input (2D image projection) of a synthetic world composed of 2D or 3D objects (see Fig. 2a for schematic). It can interact with the environment by applying forces to locations in the image, which corresponds to forces in the 2D/3D world. This architecture allows the agent to learn complex conditional rules, such as $(A\to B\land C)\land(\neg A\to D\land E)$, perform logical composition $(A\to B)\land(A\to C)\vdash A\to(B\land C)$ by learning two different rules separately but then they emerge combined during execution, and XOR operations, such as $(A\lor B)\to C)\land(A\land B\to D)$. Once the agent learns the rules, reasoning transforms into a behavioural response that changes the environment towards its internal preference. Furthermore, the agent, using the proposed architecture, shows online adaptation to unexpected changes in its environment, is robust to mild violations of its world model, and invariant to the number of objects.
@@ 4.1 Environment
We developed, Active dSprites, which is an “activated” version of the various multi-dSprites datasets that have been used in previous work on object-based visual inference (e.g. [10, 12]). Not only does it include (continuous) dynamics, but these dynamics can can be acted on by an agent (through continuous control actions that accelerate the objects). Thus, active dSprites is an interactive environment, rather than a dataset. We implemented two different scenarios in the active dSprites environment. The first, objects are 2.5-D shapes (e.g., squares, ellipses and hearts)—they have no depth dimension of their own, but can occlude each other within the depth dimension of the image. The second objects are 3D shapes (e.g., boxes, half-torus, cones) that can move in a $10\times 10\times 10$ m cubic space. These objects are much more complex to learn than the 2D version, as they have lighting and shadows, producing color gradients that need to be encoded. To randomize the environment, when an active dSprites instance is intialized, object shapes, positions, sizes and colors are all sampled uniformly at random. Initial velocities are drawn from a Normal distribution. Shape colors are sampled at discrete intervals spanning the full range of RGB-colors, while a background color is drawn from a set of evenly spaced grayscale values between black and white. Shapes are presented in random depth order. In the 3D environment the lighting sources are fixed. Code for the active dSprites environment can be found at github.com/neuro-ai-robotics/OBR.
@@ 4.2 Problem complexity analysis through baselines comparison
Figure 3: Complexity analysis through baseline comparison. Oracle algorithms have access to the objects location and dynamics, and define the upper-bound performance. RL algorithms (PPO and SAC) are tested with full observability and partial observability (RGB image pixels) (a) Algorithms performance at each frame of execution for a learnt single conditional rule $(A\to B)\land(\neg A\to C)$. Colored lines are the Mean Squared Error (MSE) between the objects location and the idealized goal location for 512 tested instances after training. Error bars are the standard deviation. (b) Average performance of the algorithms at the final frame. The lower the better.
Results are shown in Fig. 3. The left panel shows the task accuracy (Mean Squarred Error, MSE) of the current object locations and the desired state in projected image) [^7] Note that this error is over the ground truth goal as we can use the environment information to obtain the true image projection of the goal.. All methods with full observability can solve the task, having all of them similar performance. When restricting the input to RGB images (pixels) standard SAC or PPO cannot learn the conditional behavioural rule. We can see that the MSE increases along the time. OBR, while it is not as accurate as the oracle, can learn the rule. The right panel shows the final performance of the compared algorithms. Oracle algorithms show the upperbound performance as it is constrained by the dynamics of the environment.
These results first show, in line with previous works [21], that manipulating objects in the environment cannot be solved through non object-centric approaches, and highlights an important weakness of current pixel-based RL baselines. Second, it shows that a basic reasoning rule that implies interaction with the environment is already sufficiently complex. Thus, learning multiple rules and composition in this active dSprite synthetic environment is relevant.
@@ 4.2.1 Abstract behavioural reasoning with conditional rules
We evaluated the capacity of the architecture to learn proto-symbolic behavioural rules, and its generalization to different number of objects. We evaluated both 2D and 3D scenarios. Figure 4 shows an OBR agent execution for an“IfHeart” conditional $(A\to B)\land(\neg A\to C)$ and the XOR rules, and two randomized scenarios for each rule. In Fig. 4a, the first scenario, there is no heart, and boxes should be moved to the upper-left and ellipses to the middle-left. In the second one, there is a heart present, and boxes, ellipses and hearts should be moved to the top-left, center-left and bottom-left, respectively. Thus, an object’s own shape determines its goal position along the vertical axis, while the presence of a heart anywhere in the scene determines all objects’ goal positions along the horizontal axis. The agent is passive until the 5th frame. The environment row shows the visual input and the red arrows describe the forces applied. The imagined goal row, describes the output from the preference network at each frame. Note that the preference is in encoded only in the latent representation and here, for visualization purposes, we show its reconstruction using the decoder network. We let the agent to perform three frames of perceptual inference before activating the preference network to infer the current goal. The true goal is the idealized solution of this “IfHeart” task. Analogously, in Fig. 4b, the OBR agent, which has learnt the Heart XOR Square task, solves two 2d active dSprites scenarios.
@@ 4.2.2 Adaptation to changes in the environment
We evaluated the capacity of the OBR agent to recover from unexpected changes in the environment. For that purpose, we generate instances of the active dSprites that perform an object substitution in the middle of the execution. For instance, a Heart is changed by a Square during a conditional rule execution. Figure 6a describes two instances of the adaptation environment and Fig. 6b the statistical performance over 512 randomized instances. The object substitution experiment test the perception module, which needs to adapt fast to estimate the new situation, the preference network, which has to adapt the imagined goal depending on the objects that appear in the scene, and the control module that should provide the correct continuous actions to correct for the previously generated object movements. The true goal describes the idealized desired objects location after object substitution. The task error shows a decrease of performance until the system is able to recover and generate the proper behaviour to fulfill the conditional rule.
Figure 6: Adaptation. (a) Two instances of the active dSprites solved by the OBR agent which has learnt the “IfHeart” conditional rule: ((((Heart $\to$ Squares to top-right $\land$ Ellipses to middle-right)(¬$\land$($\neg$) ∧ ( ¬ Heart $\to$ Squares to top-left $\land$ Ellipses to middle-left $\land$ Hearts to down-left)))). At 4th frame one of the objects (in this case the heart is substituted by a square, forcing not only to change the object-centric perception but the conditional behaviour, as the resulting reasoning depends on the appearance of the Heart. (b) Statistical performance evaluation of 512 instances of the object substitution experiment. The blue line is the median and error bars define the inter-quartile range to deal with the skewed data due to failures.